OpenStudy (anonymous):

A Ferris wheel has a radius of 80 feet. Two particular cars are located such that the central angle between them is 165º. To the nearest tenth, what is the length of the intercepted arc between those two cars on the Ferris wheel? HELP

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

pwease?

OpenStudy (anonymous):

@poona

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

pooja help?

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

@Paki

OpenStudy (anonymous):

arc length = r*π*angle in degrees/180

OpenStudy (anonymous):

hmm

OpenStudy (jtvatsim):

@Icedragon50 Did peachpi's formula help you solve the problem?

OpenStudy (anonymous):

well i think first i have to convert 165 degrees to radians form

OpenStudy (anonymous):

then i would have \[\frac{ 11 \pi}{ 12}\]

OpenStudy (jtvatsim):

That is correct! Peachpi's formula already is doing this for you.

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

oh when i plug it into peach pi's formula i get 230.4 right?

OpenStudy (jtvatsim):

Let me check... I'm not as fast with a calculator. :)

OpenStudy (anonymous):

okie :P

OpenStudy (jtvatsim):

Yes, and that is the correct answer to the problem. Of course, don't forget units. :)

OpenStudy (anonymous):

YAY so 230.4 feet is the answer thanks so much :D

OpenStudy (jtvatsim):

You're welcome! What you did with 11pi/12 was the pi*angle in degrees/180 part of peachpi's formula. Nice work.

OpenStudy (anonymous):

sowwy but i think since peachpi gave the ormula he should get the medal :P

OpenStudy (jtvatsim):

I agree, and gave peachpi a medal as well. :)

OpenStudy (anonymous):

:D